OpenStudy (cutiecomittee123):

If sin x = 3/5 and x is in quadrant 2, what is the value of tan x/2?

OpenStudy (cutiecomittee123):

Well I think it is supposed to be solved using trig identities

OpenStudy (trojanpoem):

cosx = 1- 2 sin^2 (x/2) cosx - 1 = -2sin^2(x/2) -0.5cosx +0.5 = sin^2(x/2) cosx = -4/5 (2nd quarter) sinx/2 = sqrt(-0.5 * -4/5 + 0.5) = 3sqrt(10)/10  tanx/2 = -3sqrt(10)/sqrt(10) = -3 (2nd quarter) Any mistakes ?

OpenStudy (cutiecomittee123):

I get tan x = -3/4 since cos= -4/5 and sin=3/5 so (3/5)/(-4/5)=-3/4

OpenStudy (cutiecomittee123):

but then since its tan x/2 I think you divide that by two so tan x/2=(-3/4)/2 tanx/2=-3/8

OpenStudy (trojanpoem):

NO

OpenStudy (trojanpoem):

sin(60) = sqrt(3)/2 so sin(30) = sqrt(3)/ 4 ??? of course not , you have to use half angle identity

OpenStudy (trojanpoem):

cosx = 1- 2 sin^2 (x/2) cosx = 2cos^2(x/2) - 1 cosx = cos^2x/2 - sin^2x/2 sinx = 2 sinx/2 cosx/2

OpenStudy (trojanpoem):

tanx = 2tanx/2 / 1- tan^2x/2

OpenStudy (trojanpoem):

That's the double angle identities , you can derive the half angle ones from them.

OpenStudy (cutiecomittee123):

Okay yeah I have those in my notes. But I dont understand where and how they are used when trying to solve these kind of questions.

OpenStudy (cutiecomittee123):

??? liek I see that you plugged them in but how did you know that and when do you use this type of identity versus other ones?

OpenStudy (trojanpoem):

It's easy to notice when to use them, if it's asking for ex for sin2x and the given is sinx so you have to use this identities , sin4x and you 're given sin2x or sinx or sin0.5x it asks for sin0.5x and you have sin4x and so on.

OpenStudy (trojanpoem):

always think of them when the angles are double or half in either sin or cos or tan or cot or cosec or sec xD

OpenStudy (cutiecomittee123):

okay well it looks like tan x = 3 not -3

OpenStudy (cutiecomittee123):

I am still very confused with this math. I am going to try another question and see if it will help my understanding

OpenStudy (rccb):

im pretty sure its x=3

OpenStudy (trojanpoem):

@RCCB , x is in quadrant 2, is in the second quarter so it's negative.

OpenStudy (trojanpoem):

180 >= x >= 90 (divide the whole equation over 2) 90 >= x/2 >= 45 (first quarter ) so x/2 in the 1st , x in the 2nd

OpenStudy (rccb):

oh